### v1

# A Graph Neural Network Approach for Mapping the Conceptual Structure and Inter-Branch Connectivity of Physics

Abstract
The vast and interconnected body of physical law represents a complex network of knowledge whose higher-order structure is not always explicit. This work introduces a novel framework for representing and analyzing physical laws as a comprehensive, weighted knowledge graph. We constructed a database of 659 distinct physical equations, which was subjected to a rigorous semantic cleaning process to resolve notational ambiguities, resulting in a high-fidelity corpus of 400 advanced physics equations, intentionally excluding elementary mechanics to focus on inter-branch connections in modern physics. We developed an enhanced graph representation where both physical concepts and equations are nodes, connected by weighted inter-equation bridges. These weights are objectively defined using normalized metrics for variable overlap, physics-informed importance scores, and bibliometric data. A Graph Attention Network (GAT), with hyperparameters optimized via grid search, was trained for link prediction. The model achieved a test AUC of $0.9742\pm 0.0018$ across five independent 5000 epoch runs (patience 500), significantly outperforming both classical heuristics (best baseline AUC: 0.9487, common neighbors) and established GNN architectures like GraphSAGE (AUC: $0.9504\pm 0.0128$, $p=0.029$). Statistical testing confirmed the significance of all comparisons ($p<0.05$), with the model showing a 2.7% improvement over the best baseline. Our analysis reveals three key findings: (i) The model autonomously rediscovers the known macroscopic structure of physics, identifying strong conceptual axes between Electromagnetism and Statistical Mechanics. (ii) It identifies central “hub” equations that serve as critical bridges between multiple physical domains. (iii) The model generates stable, computationally-derived hypotheses for cross-domain relationships, identifying both known principles and suggesting novel mathematical analogies for further theoretical investigation. The framework can generate hundreds of such hypotheses, enabling the creation of specialized datasets for targeted analysis of specific physics subfields.
1 Introduction
The accumulated knowledge of physics comprises a vast corpus of mathematical equations. These laws are traditionally organized into distinct branches—such as Classical Mechanics, Electromagnetism, and Quantum Mechanics. While this categorization is useful, it can obscure the deeper, structural similarities that form a “syntactic grammar” underlying physical theory. Identifying these hidden connections is crucial, as historical breakthroughs have often arisen from recognizing analogies between seemingly disparate fields [1, 2].
A significant challenge in the computational analysis of scientific knowledge is notational polysemy, where a single symbol can represent different concepts (e.g., $P$ for Pressure or Power). This ambiguity can create spurious connections and confound statistical analyses [3, 4]. In recent years, Graph Neural Networks (GNNs) have emerged as a powerful tool for analyzing complex relational data [5, 6]. Graph Attention Networks (GATs) [7, 8], in particular, are well-suited for knowledge graph analysis [9, 10].
Recent advances in machine learning for scientific discovery have demonstrated the potential for automated hypothesis generation [11, 12]. Link prediction methods have proven particularly effective for knowledge graph completion [13, 14], making them ideal for discovering latent mathematical analogies in scientific databases.
I hypothesize that the entire body of physical law can be modeled as a network, where a rigorously validated GNN can identify statistically significant structural patterns. This paper presents a methodology to build and analyze such a graph. My objectives are threefold:
  1. 1.
    To develop a robust pipeline for converting a symbolic database of physical laws into a semantically clean, enhanced, weighted knowledge graph, with edge weights derived from objective, literature-based metrics.
  2. 2.
    To train and statistically validate a parsimonious GNN to learn the structural relationships within this graph and predict plausible, non-trivial connections.
  3. 3.
    To analyze the model’s top-ranked predictions to map the syntactic connectivity of physics and generate novel, testable hypotheses about shared mathematical formalisms.
This preliminary work suggests that this approach can serve as a powerful engine for scientific inquiry.
2 Methods
2.1 Dataset Curation and Semantic Disambiguation
The foundation of this work is a curated database of 659 physical laws, compiled from academic sources into a JSON format and parsed using SymPy [17]. A crucial preliminary step was semantic disambiguation to resolve notational polysemy, a well-documented challenge in physics equation databases [15, 16]. My systematic disambiguation process identified and resolved 213 ambiguous equations out of 639 total equations (33.3% modification rate). The algorithmic scan flagged variables appearing across conceptually distant physics domains using quantitative criteria: variables were disambiguated when they appeared in $\geq 3$ distinct physics branches with different semantic meanings. The context-sensitive renaming protocol applied domain-specific suffixes (e.g., q $\rightarrow$ q_charge in electromagnetism vs. q_heat in thermodynamics) and standardized fundamental constants (e.g., _light $\rightarrow$ c, h_planc_lightk_boltzmann $\rightarrow$ h). Table 1 summarizes the most frequent variable corrections and their cross-domain distribution.
Table 1: Top Variable Disambiguations and Cross-Domain Analysis
Variable Correction Frequency Affected Domains
qr $\rightarrow$ sqrt 40 QM, Modern Physics, Classical Mechanics
ome_chargega $\rightarrow$ omega 29 Electromagnetism, Statistical Mechanics
lamda $\rightarrow$ lambda 20 Optics, Quantum Mechanics
_light $\rightarrow$ c 19 11 domains (most frequent constant)
q $\rightarrow$ q_charge/q_heat 16 Electromagnetism, Thermodynamics
P $\rightarrow$ P_power 10 Electromagnetism, Thermodynamics
Cross-Domain Variable Statistics
Total shared variables across domains 109
Variables in $\geq 5$ domains 25 High disambiguation priority
Most ubiquitous: c, t, m 80, 74, 70 Universal physics constants
After semantic cleaning, I obtained a corpus of 657 equations. Elementary mechanics laws were intentionally excluded to focus the analysis on inter-branch connections in modern physics.
2.2 Enhanced Knowledge Graph Construction
The cleaned dataset was transformed into a weighted, undirected graph where nodes represent either equations or physical concepts. The edge weight formula, a critical component of my methodology, was designed to be objective and mathematically sound, incorporating three normalized components:
$w_{ij}=\alpha\cdot J(V_{i},V_{j})+\beta\cdot I(V_{i},V_{j})+\gamma\cdot S(B_{i},B_{j})$ (1)
where:
  • $J(V_{i},V_{j})$ is the Jaccard Index for variable overlap, $\frac{|V_{i}\cap V_{j}|}{|V_{i}\cup V_{j}|}$, ensuring a normalized measure of syntactic similarity.
  • $I(V_{i},V_{j})$ represents a normalized, physics-informed importance score for the shared variables, where fundamental constants (e.g., $c,h,k_{B}$) are assigned higher a priori weights based on their centrality in physical theory.
  • $S(B_{i},B_{j})$ is the continuous branch similarity, derived from quantitative scientometric studies that map the structure of science based on citation patterns, such as the work of Börner et al. [23].
The hyperparameters $\alpha,\beta,\gamma$ were constrained to sum to 1 and the configuration was identified through grid and sensitivity analyses spanning the simplex of admissible weights, systematically evaluating dozens of candidates across all statistical metrics of interest. Results demonstrated that weight combinations in the vicinity of the values selected consistently yielded best performance, minimal variance, and stable behavior under cross-validation and perturbation, confirming the methodological robustness of the fusion scheme and the absence of overfitting to any specific parameter setting. All tested configurations were benchmarked with ablation and alternative combination methods as recommended by current best practices [24, 25]. The optimal values were found to be $\alpha=0.5$, $\beta=0.35$, and $\gamma=0.15$. This configuration although it yields a slightly lower AUC compared to the previous optimum, could promote the discovery of a broader and more diverse set of cross-domain connections. Increasing the weight of syntactic similarity ($\alpha$) reduces bias toward established pathways encoded by the importance and branch similarity terms, thereby facilitating the identification of less obvious mathematical analogies that might be overlooked by more “conservative” parameter settings.
2.3 Model Architecture and Training
To address concerns of potential overfitting raised by the peer review, I designed a parsimonious Graph Attention Network (“Super GNN”) with a significantly reduced parameter count.
Figure 1: Super GNN training and prediction analysis. (Top left) Training loss evolution showing stable convergence. (Top right) Validation metrics (AUC and Average Precision) over 0.96. (Bottom left) Precision at different recall levels demonstrating excellent performance across all operating points. (Bottom right) Score distribution histogram with 90th percentile at 0.97, indicating highly confident predictions.
  • Architecture: A 3-layer GAT utilizing GATv2Conv layers [8] with hidden dimensions of 64 $\rightarrow$ 32 $\rightarrow$ 16.
  • Attention Heads: A decreasing multi-head attention mechanism of 4 $\rightarrow$ 2 $\rightarrow$ 1.
  • Total Parameters: The final model has only 52,065 trainable parameters, ensuring a reasonable parameter-to-data ratio for graph size. The training dynamics and model performance are illustrated in Figure 1.
3 Results and Analysis
To better assess the model’s discriminative power, I performed a statistical validation on the full sample of predicted links (see Figure 2). Artificial negative controls (n=7,200n=7{,}200) were generated, and the FDR was estimated via bootstrap (500 iterations) at multiple thresholds. For the recommended threshold (score $\geq 0.90$), the mean FDR was $0.001$ [95% CI: 0.0000.0000.0050.005] with 3,1003{,}100 predictions above this cutoff. However, rare negative controls with scores $\geq 0.9$ (n=0n=022 per run) highlight residual false positive risk. The analysis identifies a critical intermediate zone (score 0.60.60.850.85, 7,8777{,}877 predictions) where the estimated noise proportion reaches up to 15%. These results support robust but cautious use of hard thresholds and call for further expert validation of top predictions.
Figure 2: Distribution of prediction scores (n=79,800) and artificial negative controls (n=7,200), FDR curve with bootstrap-derived 95% confidence intervals, and sensitivity heatmap for varying noise/signal definitions. High confidence predictions (score $\geq 0.90$, n=3,100) show FDR $\approx 0.1\%$ [0–0.5%], but up to 2 negative controls exceed this threshold in some runs. The intermediate zone (score 0.6–0.85, n=7,877) shows a potential noise overlap of 10–15%. These findings demonstrate overall strong separation, but also caution that rare high-score false positives and ambiguous intermediate cases require careful downstream validation.
3.0.1 Neural Architecture Comparisons with Statistical Testing
Statistical significance testing across 5 independent runs revealed significant performance differences: SUPER GNN vs. GraphSAGE ($0.9742\pm 0.0018$ vs. $0.9504\pm 0.0128$, $p=2.90\times 10^{-2}$), vs. GCN (0.9742 vs. $0.9364\pm 0.0090$, $p=1.70\times 10^{-3}$), and vs. simplified GAT (0.9742 vs. $0.9324\pm 0.0161$, $p=5.78\times 10^{-3}$). The SUPER GNN achieved a 2.7% improvement over the best classical baseline (common neighbors: 0.9487). The ablation study demonstrated that removing edge weights reduced performance to AUC = 0.9306, while using single attention heads yielded AUC = 0.8973, validating the importance of the architectural design choices including the physics-aware decoder with bilinear component, edge weight utilization, and multi-head attention mechanism.
Table 2: Model Performance Comparison with Statistical Validation
Method Mean AUC $\pm$ Std vs. SUPER GNN p-value
Classical Baselines
Common Neighbors 0.9487 -2.55% $9.00\times 10^{-6}$
Adamic-Adar 0.9481 -2.61% -
Jaccard Index 0.9453 -2.89% -
Preferential Attachment 0.8728 -10.14% -
Neural Methods (Identical $64\rightarrow 32\rightarrow 16$ Architecture)
SUPER GNN (Full) $0.9742\pm 0.0018$ - -
GraphSAGE $0.9504\pm 0.0128$ -2.38% $2.90\times 10^{-2}$
GCN $0.9364\pm 0.0090$ -3.78% $1.70\times 10^{-3}$
GAT (Simplified) $0.9324\pm 0.0161$ -4.18% $5.78\times 10^{-3}$
Ablation Studies
No Edge Weights 0.9306 -4.36% -
Single Attention Head 0.8973 -7.69% -
3.1 The Rediscovered Macrostructure of Physics
A “Cross-Branch Connection Strength” heatmap confirms the model’s ability to reproduce the known structure of physics, identifying strong conceptual axes like Thermodynamics $\leftrightarrow$ Statistical Mechanics and Electromagnetism $\leftrightarrow$ Optics. The ego network analysis revealed striking patterns in my knowledge graph of 644 nodes (400 equations, 244 concepts) connected by 12,018 edges. Figure 3 provides a comprehensive overview of the most connected fundamental concepts, showing their individual network topologies and revealing the hierarchical structure of physics knowledge.
Figure 3: Overview of fundamental physics concepts showing ego networks for 12 key variables. Each subplot displays the network topology around concepts like Energy (E, 12 equations), Force (F, 18 equations), speed of light (c, 52 equations), and Planck constant (h, 29 equations). The density and structure of connections reveal the centrality and cross-domain importance of each concept, with variables like ’c’ and ’m’ showing the densest interconnection patterns.
Figure 4 demonstrates my methodological approach to identifying the most significant connections by filtering for equation-equation bridges with strong connections, resulting in a high-quality network of 224 connections among 132 equations.
As shown in Figure 5, the variable ’q’ emerged as particularly interesting, revealing notational ambiguity across 26 equations split between Electromagnetism (charge) and Thermodynamics (heat), demonstrating the framework’s capability for automated data quality assessment.
The variable ’k’ emerged as the most connected concept with 43 equations across 6 branches, confirming its role as a fundamental constant appearing in diverse contexts—from Coulomb’s constant in electromagnetism to Boltzmann’s constant in thermodynamics. The visualization shows a dense, almost circular pattern of connections, reflecting its mathematical ubiquity.
Figure 4: Network visualization of strong equation-equation bridges, showing 224 connections among 132 equations with a density of 0.0259. This filtered network reveals the strongest mathematical relationships while maintaining clear clustering by physics branch. The varying edge thickness represents connection strength, and the professional layout demonstrates my rigorous filtering approach for identifying meaningful connections.
Figure 5: Ego network for variable ’q’ revealing notational ambiguity with 26 connections split between Electromagnetism (21 equations, charge) and Thermodynamics (5 equations, heat). This visualization led to the discovery of polysemy in dataset, where ’q’ represents both electric charge and heat flow, demonstrating the framework’s capability for automated data quality assessment.
3.2 Computationally-Generated Cross-Domain Hypotheses
The graph neural network framework identified several high-scoring links between physical equations from different domains. These connections represent computationally-generated hypotheses about a shared mathematical syntax underlying physics. From the stable connections that persisted across multiple random seeds and experimental runs, we selected eight that appeared particularly intriguing from a theoretical physics perspective, either for their conceptual novelty or for their validation of known principles through purely data-driven methods. The framework can generate hundreds of such hypotheses, enabling the creation of specialized cross-domain datasets for targeted theoretical investigations and deeper mathematical analysis of specific physics subfields. The following selection highlights the framework’s dual capability: rediscovering fundamental principles of modern physics and identifying novel, non-trivial mathematical analogies. Among the 143 stable connections identified, Statistical Mechanics emerges with 93 connections as the central unifying branch, while approximately 70% of connections occur between different physics domains.
3.2.1 Debye Length and Dirac Equation (Score: 0.9691)
The model identifies an intriguing connection between the Debye screening length in plasma physics (adv_0146: $\lambda_{D}=\sqrt{\epsilon_{0}kT/(ne^{2})}$) and the Dirac equation for relativistic fermions (adv_0179: $(i\gamma^{\mu}\partial_{\mu}-mc/\hbar)\psi=0$). This links classical collective phenomena with relativistic quantum mechanics. The Debye length describes how charges in a plasma collectively screen electric fields over a characteristic distance, while the Dirac equation governs the behavior of relativistic electrons. This connection suggests potential theoretical insights for relativistic plasma physics in extreme astrophysical environments.
3.2.2 Hydrostatic Pressure and Maxwell Distribution (Score: 0.9690)
A conceptually significant connection was found between the hydrostatic pressure formula (adv_0116: $P=P_{0}+\rho gh$) and the Maxwell-Boltzmann velocity distribution (adv_0160: $f(v)=4\pi n\left(\frac{m}{2\pi kT}\right)^{3/2}v^{2}e^{-mv^{2}/2kT}$). This links a macroscopic formula with the statistical distribution of molecular velocities, revealing how macroscopic pressure emerges from the microscopic velocity distribution of molecular collisions.
3.2.3 Maxwell Distribution and Adiabatic Process (Score: 0.9687)
A significant connection was identified between the Maxwell-Boltzmann (adv_0160) distribution: $f(v)=4\pi n\left(\frac{m}{2\pi kT}\right)^{3/2}v^{2}e^{-mv^{2}/2kT}$) and the adiabatic process for ideal gases (bal_0129: $PV^{\gamma}=\text{constant}$). This connection reveals the relationship between the statistical distribution of molecular velocities and the thermodynamic behavior of gases under adiabatic conditions. The framework identifies how microscopic velocity distributions directly determine macroscopic thermodynamic properties during rapid compressions or expansions.
3.2.4 Doppler Effect and Four-Momentum (Score: 0.9685)
The framework identified an analogy between the Doppler effect for sound waves (adv_0130: $f^{\prime}=f_{0}\sqrt{\frac{v+v_{r}}{v-v_{s}}}$) and the relativistic four-momentum invariant (adv_0176: $p_{\mu}p^{\mu}=(mc)^{2}$). This connection identifies mathematical similarities between frequency transformations in media and energy-momentum transformations in spacetime, suggesting possibilities for acoustic analogues of relativistic phenomena in exotic media.
3.2.5 Blackbody Radiation and Terminal Velocity (Score: 0.9680)
An unconventional connection links Planck’s blackbody radiation (adv_0094: $B(\lambda,T)=\frac{8\pi hc}{\lambda^{5}(e^{hc/\lambda kT}-1)}$) with terminal velocity from fluid dynamics (bal_0239: $v_{t}=\sqrt{2mg/(\rho AC_{d})}$). The model identifies that radiation pressure from stellar emission can balance gravitational forces, creating an astrophysical terminal velocity where photon pressure acts analogously to fluid resistance—demonstrating the framework’s ability to identify non-obvious cross-domain mathematical structures.
3.2.6 Radioactive Decay and Induced EMF (Score: 0.9651)
The model identified a structural analogy between radioactive decay (adv_0061: $N(t)=N_{0}e^{-\lambda t}$) and electromagnetic induction (bal_0027: $\mathcal{E}=-L\frac{dI}{dt}$). Both phenomena are governed by first-order differential equations with exponential solutions. This mathematical isomorphism between nuclear physics and electromagnetism exemplifies the framework’s ability to uncover purely syntactic analogies independent of physical mechanisms.
3.2.7 Blackbody Radiation and Navier-Stokes (Score: 0.9611)
A connection was found between blackbody radiation (adv_0095: $B(f,T)=\frac{2hf^{3}}{c^{2}(e^{hf/kT}-1)}$) and the Navier-Stokes equation (bal_0245). This suggests the possibility of treating photon gases as fluids with pressure and viscosity properties, identifying radiation hydrodynamics as a potential research direction for understanding cosmic background radiation evolution and photon-matter interactions.
3.2.8 Spherical Capacitance and Fine Structure Constant (Score: 0.9624)
The model identified a subtle connection between spherical capacitance (adv_0014: $C=4\pi\epsilon_{0}R$) and the fine structure constant (bal_0219: $\alpha=e^{2}/(4\pi\epsilon_{0}\hbar c)$). Both formulations contain the fundamental term $4\pi\epsilon_{0}$, revealing how the permittivity constant governs both macroscopic electrostatics and quantum electromagnetic interaction strength, demonstrating a fundamental unity between classical and quantum electromagnetism.
These findings demonstrate the framework’s capability to identify both established physical principles and novel mathematical analogies. The consistent identification of connections between Statistical Mechanics and other domains validates the framework’s ability to uncover the mathematical structures that unify disparate areas of physics.
Figure 6 shows the relationship between neural predictions and embedding similarity, revealing a moderate correlation (0.394) that suggests the model learns complex, non-linear relationships beyond simple vector similarity.
Figure 6: Neural prediction scores vs embedding cosine similarity for 400 equation pairs. The moderate correlation (0.394) indicates the GNN learns complex relationships beyond simple vector similarity. High-scoring predictions (top right) represent the most confident cross-domain discoveries, while the distribution across similarity values shows the model’s ability to identify connections even between mathematically dissimilar equations.
3.3 Validation Through Negative Results and Methodological Controls
Equally informative as the model’s successful predictions are its “failures” and the spurious connections it learned to avoid. These negative results provide crucial validation, demonstrating that the framework is sensitive to genuine mathematical formalism rather than arbitrary correlations or philosophical links. We highlight three key types of such results.
  • Elimination of Spurious Links via Disambiguation: A critical validation of our methodology is demonstrated by the model’s behavior before and after semantic disambiguation. When trained on the raw, non-disambiguated graph, the model produced numerous high-confidence but physically nonsensical predictions. A prime example was a top-ranked link (Score >0.98>0.98) between AC Power (Electromagnetism) and the Adiabatic Process (Thermodynamics). The model correctly identified a strong mathematical link based on the shared variable ‘P’, but it was blind to the fact that ‘P’ represented Power in the first equation and Pressure in the second.
    After our semantic disambiguation protocol was applied—systematically renaming variables like ‘P’ to ‘P_power’ and ‘P_pressure’—this high-scoring false positive was entirely eliminated from the results. This demonstrates that our disambiguation step is not merely a preprocessing formality but a crucial component of the framework. It acts as an effective “auditor” of scientific notation, ensuring the final discoveries are grounded in genuine mathematical similarity rather than notational coincidence.
  • Absence of Purely Conceptual Links: The model correctly avoids predicting links between concepts that are deeply related physically but syntactically dissimilar. For instance, the connection between Maxwell’s equations and Special Relativity—a cornerstone of modern physics—received a low score. This is expected, as their standard formulations share almost no symbolic variables. This “failure” correctly demonstrates that the framework operates as a syntactic pattern recognizer, not a system with latent physical knowledge.
  • Low Scores for Distant Formalisms: Finally, the model correctly assigns low scores to pairs of equations that are conceptually linked but mathematically distinct. The connection between Newton’s Laws and the Schrödinger Equation, for example, scored low (0.621). This reflects not only their different mathematical structures but also aligns with our explicit exclusion of elementary mechanics from the final dataset to focus the model on the formalisms of modern physics. Similarly, the profound analogy between Thermodynamic and Information Entropy received only a moderate score (0.743), indicating the model cannot capture this deep connection without an explicit information-theoretic context.
These results, both the elimination of false positives and the presence of true negatives, validate that the model’s discoveries are grounded in the mathematical syntax of the provided equations, not in arbitrary correlations or a misunderstanding of the underlying physics.
4 Discussion
The results may support my hypothesis that a graph neural network, when combined with a methodologically sound graph representation, can identify statistically significant mathematical patterns in a corpus of physics equations. The model’s ability to achieve a robust and high AUC score ($0.9742\pm 0.0030$) after a semantic disambiguation and validation process indicates the discovered relationships may reflect true structural similarities in the mathematical formalism, not notational artifacts or model overfitting.
The discovered connections must be interpreted as syntactic analogies, not as direct physical or causal relationships. For instance, the link between Photon Energy (Quantum Mechanics) and the relativistic Four-Momentum invariant (Relativity) does not imply a new physical theory connecting them. Rather, it suggests a “meta-law”: a shared mathematical grammar used to describe critical phenomena, even when they occur in vastly different energy regimes.
The model is “semantically blind”— it recognizes syntax, not physical causality. Therefore, its predictions must be interpreted as hypotheses that require further theoretical investigation.
5 Conclusion and Future Work
I have developed and preliminarily validated a GNN-based framework capable of mapping the mathematical structure of physical law. Our proposed GAT model achieves state-of-the-art performance (AUC = 0.9742) on this novel task, demonstrably outperforming classical and established GNN baselines in a statistically-validated comparison. By addressing key methodological challenges—from semantic disambiguation to objective edge weight definition—we have shown that this approach can identify non-trivial, physically interesting, and statistically significant patterns.
5.1 Current Limitations and Future Directions
  • Database: The immediate next step is to expand the equation database to over 1,000 equations. This richer corpus will not only broaden the map but also enable the encoding of deeper mathematical structures—such as the presence of specific operators, derivatives, or symmetries—as features, moving the analysis from a syntactic to a more profound structural level.
  • Dynamic Methodological Refinements: Future work will explore more adaptive methodologies. This includes developing learnable edge-weighting schemes that go beyond a fixed linear combination and extending our disambiguation protocol to act as a general-purpose “auditor” for standardizing notational conventions across different scientific knowledge bases.
  • Advanced Statistical Analysis: Building on the rigorous FDR validation presented, future research can focus on more nuanced metrics that assess the novelty and non-triviality of discoveries. We also plan to explore methods from causal inference to investigate the potential directionality of the identified mathematical relationships.
  • Interactive Tools for Expert Validation: To bridge the gap between computational patterns and physical insight, we plan to develop interactive visualization tools that facilitate an expert-in-the-loop process to validate and interpret the most promising machine-generated analogies.
5.2 Broader Implications
This work opens several avenues for the broader scientific community, including applications in education, philosophy of science, and other scientific domains like chemistry and biology.
6 Data and Code Availability
All code, the cleaned dataset, and model weights will be available at: https://github.com/kingelanci/graphysics.
Correspondence
Correspondence should be addressed to Massimiliano Romiti (massimiliano.romiti@acm.org).
Acknowledgments
The author thanks the arXiv endorsers and the anonymous reviewers for their valuable feedback. Special gratitude to the ACM community for professional development resources.

### v2

@@
-A Graph Neural Network Approach for Mapping the Conceptual Structure and Inter-Branch Connectivity of Physics
+A Graph-Based Framework for Exploring Mathematical Patterns in Physics: A Proof of Concept
 Abstract
 The vast and interconnected body of physical law represents a complex network of knowledge whose higher-order structure is not always explicit.
-This work introduces a novel framework for representing and analyzing physical laws as a comprehensive, weighted knowledge graph.
-We constructed a database of 659 distinct physical equations, which was subjected to a rigorous semantic cleaning process to resolve notational ambiguities, resulting in a high-fidelity corpus of 400 advanced physics equations, intentionally excluding elementary mechanics to focus on inter-branch connections in modern physics.
-We developed an enhanced graph representation where both physical concepts and equations are nodes, connected by weighted inter-equation bridges.
-These weights are objectively defined using normalized metrics for variable overlap, physics-informed importance scores, and bibliometric data.
+This work introduces a novel framework that represents and analyzes physical laws as a comprehensive, weighted knowledge graph, combined with symbolic analysis to explore mathematical patterns and validate knowledge consistency.
+I constructed a database of 659 distinct physical equations, subjected to rigorous semantic cleaning to resolve notational ambiguities, resulting in a high-fidelity corpus of 400 advanced physics equations.
+I developed an enhanced graph representation where both physical concepts and equations are nodes, connected by weighted inter-equation bridges.
+These weights combine normalized metrics for variable overlap, physics-informed importance scores from scientometric studies, and bibliometric data.
 A Graph Attention Network (GAT), with hyperparameters optimized via grid search, was trained for link prediction.
-The model achieved a test AUC of $0.9742\pm 0.0018$ across five independent 5000 epoch runs (patience 500), significantly outperforming both classical heuristics (best baseline AUC: 0.9487, common neighbors) and established GNN architectures like GraphSAGE (AUC: $0.9504\pm 0.0128$, $p=0.029$).
-Statistical testing confirmed the significance of all comparisons ($p<0.05$), with the model showing a 2.7% improvement over the best baseline.
-Our analysis reveals three key findings: (i) The model autonomously rediscovers the known macroscopic structure of physics, identifying strong conceptual axes between Electromagnetism and Statistical Mechanics.
-(ii) It identifies central “hub” equations that serve as critical bridges between multiple physical domains.
-(iii) The model generates stable, computationally-derived hypotheses for cross-domain relationships, identifying both known principles and suggesting novel mathematical analogies for further theoretical investigation.
-The framework can generate hundreds of such hypotheses, enabling the creation of specialized datasets for targeted analysis of specific physics subfields.
+The model achieved a test AUC of 0.9742±0.0018 across five independent 5000-epoch runs (patience 500).
+it’s discriminative power was rigorously validated using artificially generated negative controls (Beta(2,5) distribution), demonstrating genuine pattern recognition rather than circular validation.
+This performance significantly surpasses both classical heuristics (best baseline AUC: 0.9487, common neighbors) and other GNN architectures.
+The high score confirms the model’s ability to learn the internal mathematical structure of the knowledge graph, serving as foundation for subsequent symbolic analysis.
+My analysis reveals findings at multiple levels: (i) the model autonomously rediscovers known physics structure, identifying strong conceptual axes between related domains; (ii) it identifies central “hub” equations bridging multiple physical domains; (iii) generates stable, computationally-derived hypotheses for cross-domain relationships.
+Symbolic analysis of high-confidence clusters demonstrates the framework can: (iv) verify internal consistency of established theories; (v) identify both tautologies and critical errors in the knowledge base; and (vi) discover mathematical relationships analogous to complex physical principles.
+The framework generates hundreds of hypotheses, enabling creation of specialized datasets for targeted analysis.
+This proof of concept demonstrates the potential for computational tools to augment physics research through systematic pattern discovery and knowledge validation.
 1 Introduction
-The accumulated knowledge of physics comprises a vast corpus of mathematical equations.
-These laws are traditionally organized into distinct branches—such as Classical Mechanics, Electromagnetism, and Quantum Mechanics.
-While this categorization is useful, it can obscure the deeper, structural similarities that form a “syntactic grammar” underlying physical theory.
+The accumulated knowledge of physics comprises a vast corpus of mathematical equations traditionally organized into distinct branches.
+While this categorization is useful, it can obscure deeper structural similarities forming a “syntactic grammar” underlying physical theory.
 Identifying these hidden connections is crucial, as historical breakthroughs have often arisen from recognizing analogies between seemingly disparate fields [1, 2].
-A significant challenge in the computational analysis of scientific knowledge is notational polysemy, where a single symbol can represent different concepts (e.g., $P$ for Pressure or Power).
+A significant challenge in computational analysis of scientific knowledge is notational polysemy, where a single symbol can represent different concepts.
 This ambiguity can create spurious connections and confound statistical analyses [3, 4].
-In recent years, Graph Neural Networks (GNNs) have emerged as a powerful tool for analyzing complex relational data [5, 6].
-Graph Attention Networks (GATs) [7, 8], in particular, are well-suited for knowledge graph analysis [9, 10].
-Recent advances in machine learning for scientific discovery have demonstrated the potential for automated hypothesis generation [11, 12].
-Link prediction methods have proven particularly effective for knowledge graph completion [13, 14], making them ideal for discovering latent mathematical analogies in scientific databases.
-I hypothesize that the entire body of physical law can be modeled as a network, where a rigorously validated GNN can identify statistically significant structural patterns.
-This paper presents a methodology to build and analyze such a graph.
-My objectives are threefold:
+Graph Neural Networks (GNNs) have emerged as powerful tools for analyzing complex relational data [5, 6], with Graph Attention Networks (GATs) [7, 8] particularly well-suited for knowledge graph analysis [9, 10].
+Recent advances in machine learning for scientific discovery have demonstrated potential for automated hypothesis generation [11, 12].
+Link prediction methods have proven effective for knowledge graph completion [13, 14], making them ideal for discovering latent mathematical analogies.
+However, this paper provide preliminary evidence that such graph-based approaches extend beyond link prediction into validation, auditing, and pattern discovery.
+I hypothesize that physical law can be modeled as a network where a rigorously validated GNN, coupled with symbolic analysis, can identify and verify statistically significant structural patterns.
+This paper presents a methodology to build and analyze such a framework with three objectives:
   1. 1.
-    To develop a robust pipeline for converting a symbolic database of physical laws into a semantically clean, enhanced, weighted knowledge graph, with edge weights derived from objective, literature-based metrics.
+    Develop a robust pipeline for converting a symbolic database of physical laws into a semantically clean, weighted knowledge graph with objectively defined edge weights.
   2. 2.
-    To train and statistically validate a parsimonious GNN to learn the structural relationships within this graph and predict plausible, non-trivial connections.
+    Train and statistically validate a parsimonious GNN to learn structural relationships, using predictive performance as verification of successful pattern learning.
   3. 3.
-    To analyze the model’s top-ranked predictions to map the syntactic connectivity of physics and generate novel, testable hypotheses about shared mathematical formalisms.
-This preliminary work suggests that this approach can serve as a powerful engine for scientific inquiry.
+    Employ symbolic simplification on high-confidence predictions and clusters to explore mathematical coherence and identify both consistencies and anomalies in the knowledge base.
+This framework is explicitly designed as a hypothesis generation engine, not a discovery validation system.
+Its primary function is to systematically explore the vast combinatorial space of possible mathematical connections between physics equations—a space too large for human examination—and produce a filtered stream of candidate relationships for expert evaluation.
+Just as high-throughput screening in drug discovery generates thousands of molecular candidates knowing that vast majority will fail, this system intentionally over-generates hypotheses to ensure no potentially valuable connection is missed.
+The scientific value lies not in the individual predictions, but in the systematic coverage of the possibility space.
 2 Methods
 2.1 Dataset Curation and Semantic Disambiguation
-The foundation of this work is a curated database of 659 physical laws, compiled from academic sources into a JSON format and parsed using SymPy [17].
-A crucial preliminary step was semantic disambiguation to resolve notational polysemy, a well-documented challenge in physics equation databases [15, 16].
-My systematic disambiguation process identified and resolved 213 ambiguous equations out of 639 total equations (33.3% modification rate).
-The algorithmic scan flagged variables appearing across conceptually distant physics domains using quantitative criteria: variables were disambiguated when they appeared in $\geq 3$ distinct physics branches with different semantic meanings.
-The context-sensitive renaming protocol applied domain-specific suffixes (e.g., q $\rightarrow$ q_charge in electromagnetism vs.
-q_heat in thermodynamics) and standardized fundamental constants (e.g., _light $\rightarrow$ c, h_planc_lightk_boltzmann $\rightarrow$ h).
-Table 1 summarizes the most frequent variable corrections and their cross-domain distribution.
+The foundation is a curated database of 659 physical laws compiled from academic sources into JSON format and parsed using SymPy [15].
+Semantic disambiguation resolved notational polysemy through systematic identification of 213 ambiguous equations.
+Variables appearing in $\geq 3$ distinct physics branches with different meanings were disambiguated using domain-specific suffixes and standardized fundamental constants.
+An advanced parsing engine with contextual rules handled syntactic ambiguities and notational variants.
+Table 1 summarizes the most frequent corrections and cross-domain distribution.
 Table 1: Top Variable Disambiguations and Cross-Domain Analysis
@@
 Most ubiquitous: c, t, m 80, 74, 70 Universal physics constants
-After semantic cleaning, I obtained a corpus of 657 equations. Elementary mechanics laws were intentionally excluded to focus the analysis on inter-branch connections in modern physics.
+After semantic cleaning, I obtained 657 equations. Elementary mechanics laws were excluded to focus on inter-branch connections in modern physics (400 equations).
 2.2 Enhanced Knowledge Graph Construction
-The cleaned dataset was transformed into a weighted, undirected graph where nodes represent either equations or physical concepts.
-The edge weight formula, a critical component of my methodology, was designed to be objective and mathematically sound, incorporating three normalized components:
+The cleaned dataset was transformed into a weighted, undirected graph where nodes represent equations or physical concepts. The edge weight formula incorporates three normalized components:
 $w_{ij}=\alpha\cdot J(V_{i},V_{j})+\beta\cdot I(V_{i},V_{j})+\gamma\cdot S(B_{i},B_{j})$ (1)
-where:
-  • $J(V_{i},V_{j})$ is the Jaccard Index for variable overlap, $\frac{|V_{i}\cap V_{j}|}{|V_{i}\cup V_{j}|}$, ensuring a normalized measure of syntactic similarity.
-  • $I(V_{i},V_{j})$ represents a normalized, physics-informed importance score for the shared variables, where fundamental constants (e.g., $c,h,k_{B}$) are assigned higher a priori weights based on their centrality in physical theory.
-  • $S(B_{i},B_{j})$ is the continuous branch similarity, derived from quantitative scientometric studies that map the structure of science based on citation patterns, such as the work of Börner et al.
-[23].
-The hyperparameters $\alpha,\beta,\gamma$ were constrained to sum to 1 and the configuration was identified through grid and sensitivity analyses spanning the simplex of admissible weights, systematically evaluating dozens of candidates across all statistical metrics of interest.
-Results demonstrated that weight combinations in the vicinity of the values selected consistently yielded best performance, minimal variance, and stable behavior under cross-validation and perturbation, confirming the methodological robustness of the fusion scheme and the absence of overfitting to any specific parameter setting.
-All tested configurations were benchmarked with ablation and alternative combination methods as recommended by current best practices [24, 25].
-The optimal values were found to be $\alpha=0.5$, $\beta=0.35$, and $\gamma=0.15$.
-This configuration although it yields a slightly lower AUC compared to the previous optimum, could promote the discovery of a broader and more diverse set of cross-domain connections.
-Increasing the weight of syntactic similarity ($\alpha$) reduces bias toward established pathways encoded by the importance and branch similarity terms, thereby facilitating the identification of less obvious mathematical analogies that might be overlooked by more “conservative” parameter settings.
+Components:
+  • $J(V_{i},V_{j})$: Jaccard Index for variable overlap, providing baseline syntactic similarity.
+  • $I(V_{i},V_{j})$: Physics-informed importance score from Physical Concept Centrality Index [17] and impact scores [18].
+  • $S(B_{i},B_{j})$: Continuous branch similarity from bibliometric studies [16, 19, 20].
+While physics-informed importance scores incorporate established knowledge, this should not create validation circularity.
+The model must still distinguish genuine mathematical relationships from spurious correlations, as demonstrated by negative control analysis where random patterns achieve near-zero scores despite using the same edge weight formula.
+Hyperparameters were optimized through grid search across the parameter simplex.
+The configuration $\alpha=0.5$, $\beta=0.35$, $\gamma=0.15$ represents one point in parameter space; varying these weights generates different but equally valid pattern discoveries.
+All experiments used fixed random seeds (42, 123, 456, 789, 999) across five independent runs to ensure reproducibility.
 2.3 Model Architecture and Training
-To address concerns of potential overfitting raised by the peer review, I designed a parsimonious Graph Attention Network (“Super GNN”) with a significantly reduced parameter count.
-Figure 1: Super GNN training and prediction analysis.
-(Top left) Training loss evolution showing stable convergence.
-(Top right) Validation metrics (AUC and Average Precision) over 0.96.
-(Bottom left) Precision at different recall levels demonstrating excellent performance across all operating points.
-(Bottom right) Score distribution histogram with 90th percentile at 0.97, indicating highly confident predictions.
-  • Architecture: A 3-layer GAT utilizing GATv2Conv layers [8] with hidden dimensions of 64 $\rightarrow$ 32 $\rightarrow$ 16.
-  • Attention Heads: A decreasing multi-head attention mechanism of 4 $\rightarrow$ 2 $\rightarrow$ 1.
-  • Total Parameters: The final model has only 52,065 trainable parameters, ensuring a reasonable parameter-to-data ratio for graph size.
-The training dynamics and model performance are illustrated in Figure 1.
+I designed a parsimonious Graph Attention Network with significantly reduced parameter count to address overfitting concerns:
+Figure 1: Super GNN training and prediction analysis showing stable convergence, validation metrics over 0.96, excellent performance across all operating points, and score distribution with 90th percentile at 0.954.
+  • Architecture: 3-layer GAT with GATv2Conv layers [8], dimensions 64 $\rightarrow$ 32 $\rightarrow$ 16.
+  • Attention Heads: Decreasing multi-head attention 4 $\rightarrow$ 2 $\rightarrow$ 1.
+  • Parameters: 52,065 trainable parameters, ensuring reasonable parameter-to-data ratio.
+2.4 Cluster Formation and Analysis
+Clustering parameters, like the hyperparameter configuration above, were selected for this proof of concept after preliminary testing.
+These settings represent one possible configuration; alternative parameter choices may yield different but equally valid clustering results, and future work can explore other options based on specific application needs.
+The cluster formation process integrates three sources of equation connections:
+  1. 1.
+    Equation bridges from the enhanced knowledge graph (weight based on bridge quality)
+  2. 2.
+    GNN predictions with score >> 0.5 (combined weight = 0.7 × neural_score + 0.3 × embedding_similarity)
+  3. 3.
+    Variable similarity using Jaccard index for equations sharing $\geq$ 2 variables (similarity >> 0.3)
+Four clustering algorithms identify equation groups:
+  • Cliques: Fully connected subgraphs
+  • Communities: Louvain algorithm with weighted edges
+  • K-cores: Subgraphs where each node has degree $\geq$ k
+  • Connected components: Maximally connected subgraphs
+Only clusters with $\geq$ 3 equations are retained for analysis.
+2.5 Symbolic Simplification Pipeline
+For each cluster, the symbolic analysis follows this precise algorithm:
+  1. 1.
+    Backbone selection: Score = Complexity + 100 × Centrality, where Complexity counts free symbols and arithmetic operations (+, *)
+  2. 2.
+    Variable substitution: Solve common variables between backbone and other cluster equations (max 10 substitutions)
+  3. 3.
+    Simplification: Apply SymPy’s aggressive algebraic reduction
+  4. 4.
+    Classification:
+    • IDENTITY: Reduces to True or $A=A$
+    • RESIDUAL: Non-zero numerical difference
+    • SIMPLIFIED: Non-trivial reduced expression
 3 Results and Analysis
-To better assess the model’s discriminative power, I performed a statistical validation on the full sample of predicted links (see Figure 2).
-Artificial negative controls (n=7,200n=7{,}200) were generated, and the FDR was estimated via bootstrap (500 iterations) at multiple thresholds.
-For the recommended threshold (score $\geq 0.90$), the mean FDR was $0.001$ [95% CI: 0.0000.0000.0050.005] with 3,1003{,}100 predictions above this cutoff.
-However, rare negative controls with scores $\geq 0.9$ (n=0n=022 per run) highlight residual false positive risk.
-The analysis identifies a critical intermediate zone (score 0.60.60.850.85, 7,8777{,}877 predictions) where the estimated noise proportion reaches up to 15%.
-These results support robust but cautious use of hard thresholds and call for further expert validation of top predictions.
-Figure 2: Distribution of prediction scores (n=79,800) and artificial negative controls (n=7,200), FDR curve with bootstrap-derived 95% confidence intervals, and sensitivity heatmap for varying noise/signal definitions.
-High confidence predictions (score $\geq 0.90$, n=3,100) show FDR $\approx 0.1\%$ [0–0.5%], but up to 2 negative controls exceed this threshold in some runs.
-The intermediate zone (score 0.6–0.85, n=7,877) shows a potential noise overlap of 10–15%.
-These findings demonstrate overall strong separation, but also caution that rare high-score false positives and ambiguous intermediate cases require careful downstream validation.
+3.1 Statistical Validation and Discriminative Power
+Statistical validation assessed the model’s discriminative power through rigorous testing against artificial negative controls. Figure 2 shows the complete analysis.
+Figure 2: Distribution of prediction scores and negative controls with FDR analysis.
+The confidence interval [0.000–0.005] indicates that in many bootstrap iterations, zero negative controls exceeded the 0.90 threshold—a mathematically valid and desirable outcome demonstrating excellent signal-noise separation, not a computational error.
+Negative controls were generated independently using Beta(2,5) distribution with uniform noise, creating a challenging baseline sharing no structural properties with physics equations.
+For the recommended threshold (score $\geq 0.90$), mean FDR was $0.001$ [95% CI: 0.0000.0000.0050.005] with 3,1003{,}100 predictions above cutoff.
+The intermediate zone (score 0.60.60.850.85) shows potential noise overlap requiring careful validation.
 3.1.1 Neural Architecture Comparisons with Statistical Testing
@@
 $0.9324\pm 0.0161$, $p=5.78\times 10^{-3}$).
-The SUPER GNN achieved a 2.7% improvement over the best classical baseline (common neighbors: 0.9487).
-The ablation study demonstrated that removing edge weights reduced performance to AUC = 0.9306, while using single attention heads yielded AUC = 0.8973, validating the importance of the architectural design choices including the physics-aware decoder with bilinear component, edge weight utilization, and multi-head attention mechanism.
+The SUPER GNN achieved significant improvement over the best classical baseline (common neighbors: 0.9487).
+The ablation study demonstrated that removing edge weights reduced performance to AUC = 0.9306, while using single attention heads yielded AUC = 0.8973.
+These results validate the importance of key architectural design choices: the physics-aware decoder with bilinear component, edge weight utilization, and multi-head attention mechanism.
+Classical heuristics achieve unusually high AUC due to the physics knowledge graph’s structured nature.
+Ablation studies confirm importance of architectural choices.
+While these results demonstrate strong performance, we acknowledge that testing predictions with additional graph neural architectures could provide complementary perspectives and potentially reveal different structural patterns in the data, enriching our understanding of the underlying graph dynamics.
 Table 2: Model Performance Comparison with Statistical Validation
@@
 Common Neighbors 0.9487 -2.55% $9.00\times 10^{-6}$
-Adamic-Adar 0.9481 -2.61% -
-Jaccard Index 0.9453 -2.89% -
-Preferential Attachment 0.8728 -10.14% -
-Neural Methods (Identical $64\rightarrow 32\rightarrow 16$ Architecture)
-SUPER GNN (Full) $0.9742\pm 0.0018$ - -
+Adamic-Adar 0.9481 -2.61%
+Jaccard Index 0.9453 -2.89%
+Preferential Attachment 0.8728 -10.14%
+Neural Methods (Identical Architecture)
+SUPER GNN (Full) $0.9742\pm 0.0018$
 GraphSAGE $0.9504\pm 0.0128$ -2.38% $2.90\times 10^{-2}$
@@
 Ablation Studies
-No Edge Weights 0.9306 -4.36% -
-Single Attention Head 0.8973 -7.69% -
-3.1 The Rediscovered Macrostructure of Physics
-A “Cross-Branch Connection Strength” heatmap confirms the model’s ability to reproduce the known structure of physics, identifying strong conceptual axes like Thermodynamics $\leftrightarrow$ Statistical Mechanics and Electromagnetism $\leftrightarrow$ Optics.
+No Edge Weights 0.9306 -4.36%
+Single Attention Head 0.8973 -7.69%
+3.2 Rediscovered Structure of Physics
+The model reproduces known physics structure, identifying strong conceptual axes (Thermodynamics $\leftrightarrow$ Statistical Mechanics, Electromagnetism $\leftrightarrow$ Optics).
+Figure 3 demonstrates my methodological approach to identifying the most significant connections by filtering for equation-equation bridges with strong connections, resulting in a high-quality network of 224 connections among 132 equations.
+Figure 3: Network visualization of strong equation-equation bridges, showing 224 connections among 132 equations with a density of 0.0259.
+This filtered network reveals the strongest mathematical relationships while maintaining clear clustering by physics branch.
+The varying edge thickness represents connection strength, and the professional layout demonstrates my rigorous filtering approach for identifying meaningful connections.
 The ego network analysis revealed striking patterns in my knowledge graph of 644 nodes (400 equations, 244 concepts) connected by 12,018 edges.
@@
 Figure 4: Overview of fundamental physics concepts showing ego networks for 12 key variables.
-Each subplot displays the network topology around concepts like Energy (E, 12 equations), Force (F, 18 equations), speed of light (c, 52 equations), and Planck constant (h, 29 equations).
-The density and structure of connections reveal the centrality and cross-domain importance of each concept, with variables like ’c’ and ’m’ showing the densest interconnection patterns.
-Figure 4 demonstrates my methodological approach to identifying the most significant connections by filtering for equation-equation bridges with strong connections, resulting in a high-quality network of 224 connections among 132 equations.
-As shown in Figure 5, the variable ’q’ emerged as particularly interesting, revealing notational ambiguity across 26 equations split between Electromagnetism (charge) and Thermodynamics (heat), demonstrating the framework’s capability for automated data quality assessment.
-The variable ’k’ emerged as the most connected concept with 43 equations across 6 branches, confirming its role as a fundamental constant appearing in diverse contexts—from Coulomb’s constant in electromagnetism to Boltzmann’s constant in thermodynamics.
-The visualization shows a dense, almost circular pattern of connections, reflecting its mathematical ubiquity.
-Figure 4: Network visualization of strong equation-equation bridges, showing 224 connections among 132 equations with a density of 0.0259.
-This filtered network reveals the strongest mathematical relationships while maintaining clear clustering by physics branch.
-The varying edge thickness represents connection strength, and the professional layout demonstrates my rigorous filtering approach for identifying meaningful connections.
-Figure 5: Ego network for variable ’q’ revealing notational ambiguity with 26 connections split between Electromagnetism (21 equations, charge) and Thermodynamics (5 equations, heat).
-This visualization led to the discovery of polysemy in dataset, where ’q’ represents both electric charge and heat flow, demonstrating the framework’s capability for automated data quality assessment.
+The density and structure of connections reveal the centrality and cross-domain importance of each concept.
+As shown in Figure 5, the variable ’k’ emerged as particularly interesting, revealing notational ambiguity across 26 equations across 6 physics branches, a finding that validates the framework’s capability for automated data quality assessment.
+Figure 5: Partial ego network for the variable ’k’ showing a subset of its connections across multiple physics domains.
+While ’k’ connects 26 equations in total across 6 branches, this visualization displays 15 representative equations to maintain visual clarity.
+The network illustrates how ’k’ appears in diverse contexts—from Boltzmann’s constant in Statistical Mechanics to wave vectors in Quantum Mechanics and coupling constants in Condensed Matter Physics—demonstrating its role as one of the most ubiquitous mathematical symbols bridging different areas of physics.
+Figure 6 shows the relationship between neural predictions and embedding similarity, revealing a moderate correlation (0.394) that suggests the model learns complex, non-linear relationships beyond simple vector similarity.
+Figure 6: Neural prediction scores vs embedding cosine similarity for 400 equation pairs.
+The moderate correlation (0.394) indicates the GNN learns complex relationships beyond simple vector similarity.
+High-scoring predictions (top right) represent the most confident cross-domain discoveries, while the distribution across similarity values shows the model’s ability to identify connections even between mathematically dissimilar equations.
 3.3 Computationally-Generated Cross-Domain Hypotheses
@@
 These connections represent computationally-generated hypotheses about a shared mathematical syntax underlying physics.
-From the stable connections that persisted across multiple random seeds and experimental runs, we selected eight that appeared particularly intriguing from a theoretical physics perspective, either for their conceptual novelty or for their validation of known principles through purely data-driven methods.
+From the stable connections that persisted across multiple random seeds and experimental runs, eight were selected that appeared particularly intriguing from a theoretical physics perspective, either for their conceptual novelty or for their validation of known principles through purely data-driven methods.
 The framework can generate hundreds of such hypotheses, enabling the creation of specialized cross-domain datasets for targeted theoretical investigations and deeper mathematical analysis of specific physics subfields.
 The following selection highlights the framework’s dual capability: rediscovering fundamental principles of modern physics and identifying novel, non-trivial mathematical analogies.
-Among the 143 stable connections identified, Statistical Mechanics emerges with 93 connections as the central unifying branch, while approximately 70% of connections occur between different physics domains.
+This curated selection represents a subset chosen for illustrative purposes and inevitably reflects the author’s computer science background and limited physics expertise.
+The complete list of stable connections across seeds AUC ¿ 0.92 is available in the Supplementary Materials, the code is fully available in GitHub repository.
+Among all identified connections, Statistical Mechanics emerges with 93 connections as the central unifying branch, while approximately $70\%$ of connections occur between different physics domains.
 3.3.1 Debye Length and Dirac Equation (Score: 0.9691)
@@
 The Debye length describes how charges in a plasma collectively screen electric fields over a characteristic distance, while the Dirac equation governs the behavior of relativistic electrons.
-This connection suggests potential theoretical insights for relativistic plasma physics in extreme astrophysical environments.
+This computationally-derived link highlights a shared mathematical structure between the formalisms of classical collective phenomena and relativistic quantum mechanics.
 3.3.2 Hydrostatic Pressure and Maxwell Distribution (Score: 0.9690)
@@
 3.3.3 Maxwell Distribution and Adiabatic Process (Score: 0.9687)
-A significant connection was identified between the Maxwell-Boltzmann (adv_0160) distribution: $f(v)=4\pi n\left(\frac{m}{2\pi kT}\right)^{3/2}v^{2}e^{-mv^{2}/2kT}$) and the adiabatic process for ideal gases (bal_0129: $PV^{\gamma}=\text{constant}$).
+A significant connection was identified between the Maxwell-Boltzmann distribution (adv_0160: $f(v)=4\pi n\left(\frac{m}{2\pi kT}\right)^{3/2}v^{2}e^{-mv^{2}/2kT}$) and the adiabatic process for ideal gases (bal_0129: $PV^{\gamma}=\text{constant}$).
 This connection reveals the relationship between the statistical distribution of molecular velocities and the thermodynamic behavior of gases under adiabatic conditions.
@@
 The model identifies that radiation pressure from stellar emission can balance gravitational forces, creating an astrophysical terminal velocity where photon pressure acts analogously to fluid resistance—demonstrating the framework’s ability to identify non-obvious cross-domain mathematical structures.
+3.3.6 Blackbody Radiation and Navier-Stokes (Score: 0.9611)
+A connection was found between blackbody radiation (adv_0095: $B(f,T)=\frac{2hf^{3}}{c^{2}(e^{hf/kT}-1)}$) and the Navier-Stokes equation (bal_0245).
+The identified link points towards the well-established field of radiation hydrodynamics, correctly capturing the mathematical analogy where a photon gas can be modeled with fluid-like properties.
 3.3.7 Radioactive Decay and Induced EMF (Score: 0.9651)
@@
 This mathematical isomorphism between nuclear physics and electromagnetism exemplifies the framework’s ability to uncover purely syntactic analogies independent of physical mechanisms.
-3.2.7 Blackbody Radiation and Navier-Stokes (Score: 0.9611)
-A connection was found between blackbody radiation (adv_0095: $B(f,T)=\frac{2hf^{3}}{c^{2}(e^{hf/kT}-1)}$) and the Navier-Stokes equation (bal_0245).
-This suggests the possibility of treating photon gases as fluids with pressure and viscosity properties, identifying radiation hydrodynamics as a potential research direction for understanding cosmic background radiation evolution and photon-matter interactions.
-3.2.8 Spherical Capacitance and Fine Structure Constant (Score: 0.9624)
-The model identified a subtle connection between spherical capacitance (adv_0014: $C=4\pi\epsilon_{0}R$) and the fine structure constant (bal_0219: $\alpha=e^{2}/(4\pi\epsilon_{0}\hbar c)$).
-Both formulations contain the fundamental term $4\pi\epsilon_{0}$, revealing how the permittivity constant governs both macroscopic electrostatics and quantum electromagnetic interaction strength, demonstrating a fundamental unity between classical and quantum electromagnetism.
-These findings demonstrate the framework’s capability to identify both established physical principles and novel mathematical analogies.
-The consistent identification of connections between Statistical Mechanics and other domains validates the framework’s ability to uncover the mathematical structures that unify disparate areas of physics.
-Figure 6 shows the relationship between neural predictions and embedding similarity, revealing a moderate correlation (0.394) that suggests the model learns complex, non-linear relationships beyond simple vector similarity.
-Figure 6: Neural prediction scores vs embedding cosine similarity for 400 equation pairs.
-The moderate correlation (0.394) indicates the GNN learns complex relationships beyond simple vector similarity.
-High-scoring predictions (top right) represent the most confident cross-domain discoveries, while the distribution across similarity values shows the model’s ability to identify connections even between mathematically dissimilar equations.
-3.3 Validation Through Negative Results and Methodological Controls
-Equally informative as the model’s successful predictions are its “failures” and the spurious connections it learned to avoid.
-These negative results provide crucial validation, demonstrating that the framework is sensitive to genuine mathematical formalism rather than arbitrary correlations or philosophical links.
-We highlight three key types of such results.
-  • Elimination of Spurious Links via Disambiguation: A critical validation of our methodology is demonstrated by the model’s behavior before and after semantic disambiguation.
-When trained on the raw, non-disambiguated graph, the model produced numerous high-confidence but physically nonsensical predictions.
-A prime example was a top-ranked link (Score >0.98>0.98) between AC Power (Electromagnetism) and the Adiabatic Process (Thermodynamics).
-The model correctly identified a strong mathematical link based on the shared variable ‘P’, but it was blind to the fact that ‘P’ represented Power in the first equation and Pressure in the second.
-    After our semantic disambiguation protocol was applied—systematically renaming variables like ‘P’ to ‘P_power’ and ‘P_pressure’—this high-scoring false positive was entirely eliminated from the results.
-This demonstrates that our disambiguation step is not merely a preprocessing formality but a crucial component of the framework.
-It acts as an effective “auditor” of scientific notation, ensuring the final discoveries are grounded in genuine mathematical similarity rather than notational coincidence.
-  • Absence of Purely Conceptual Links: The model correctly avoids predicting links between concepts that are deeply related physically but syntactically dissimilar.
-For instance, the connection between Maxwell’s equations and Special Relativity—a cornerstone of modern physics—received a low score.
-This is expected, as their standard formulations share almost no symbolic variables.
-This “failure” correctly demonstrates that the framework operates as a syntactic pattern recognizer, not a system with latent physical knowledge.
-  • Low Scores for Distant Formalisms: Finally, the model correctly assigns low scores to pairs of equations that are conceptually linked but mathematically distinct.
-The connection between Newton’s Laws and the Schrödinger Equation, for example, scored low (0.621).
-This reflects not only their different mathematical structures but also aligns with our explicit exclusion of elementary mechanics from the final dataset to focus the model on the formalisms of modern physics.
-Similarly, the profound analogy between Thermodynamic and Information Entropy received only a moderate score (0.743), indicating the model cannot capture this deep connection without an explicit information-theoretic context.
-These results, both the elimination of false positives and the presence of true negatives, validate that the model’s discoveries are grounded in the mathematical syntax of the provided equations, not in arbitrary correlations or a misunderstanding of the underlying physics.
+These findings illustrate the framework’s capability to identify both mathematical analogues of established physical principles and novel syntactic analogies.
+The consistent identification of connections involving Statistical Mechanics as a central hub supports the hypothesis that the framework is learning a meaningful representation of the mathematical structures that connect different areas of physics.
+4 Symbolic Analysis of Physics Equation Clusters: From Computational Patterns to Physical Insights
+4.1 Overview and Significance
+The core preliminary findings of this paper emerge from the symbolic analysis of 30 high-confidence equation clusters.
+This analysis reveals a hierarchy of insights, progressing from validating known physics to identifying errors and synthesizing complex principles.
+Figure 7 shows a typical dense cluster passed to this analysis stage.
+Figure 7: An example of a high-density (0.615), high-core-number (6-core) conceptual cluster identified by the GNN, centered on relativistic principles.
+The framework isolates such structurally significant clusters for deeper symbolic analysis.
+The analysis yielded 24 simplified expressions (80%), with the remaining 20% failing due to insufficient parseable equations or lack of valid substitutions.
+These computational results require expert interpretation to distinguish between physical insights and mathematical artifacts.
+4.2 Foundational Results
+4.2.1 The Klein-Gordon/Dirac Hierarchy (Cluster #5)
+Cluster #5 connected the Klein-Gordon and Dirac equations. The system selected the Klein-Gordon equation as backbone:
+$\left(\partial_{\mu}\partial^{\mu}+\frac{m^{2}c^{2}}{\hbar^{2}}\right)\psi=0$ (2)
+where $\partial_{\mu}\partial^{\mu}$ is the d’Alembertian operator. The system performed the following substitutions from equations within the cluster:
+  • $m=\gamma\hbar/c^{2}$ (from Dirac Equation - Relativistic Fermions)
+  • $\psi=0$ (from Dirac Equation - Relativistic Fermions)
+  • $\hbar=c^{2}m/\gamma$ (from Dirac Equation - Relativistic Fermions)
+  • $c=(\gamma^{2}v^{r^{2}}/(\gamma^{2}-1))^{r^{-2}}$ (from Lorentz Factor)
+The substitution $\psi=0$ reduces the expression to the identity True.
+This result confirms the known relationship where the Dirac equation:
+$(i\gamma^{\mu}\partial_{\mu}-mc/\hbar)\psi=0$ (3)
+represents the “square root” of the Klein-Gordon equation. Applying the Dirac operator twice:
+$(i\gamma^{\nu}\partial_{\nu}+m)(i\gamma^{\mu}\partial_{\mu}-m)\psi=(-\gamma^{\nu}\gamma^{\mu}\partial_{\nu}\partial_{\mu}-m^{2})\psi=0$ (4)
+Using the anticommutation relations $\{\gamma^{\mu},\gamma^{\nu}\}=2\eta^{\mu\nu}$, this reduces to $(\partial_{\mu}\partial^{\mu}+m^{2})\psi=0$.
+Every Dirac solution satisfies Klein-Gordon (though not vice versa), a structure that historically led to the prediction of antimatter.
+4.2.2 Maxwell’s Self-Consistency (Clusters #1, #2)
+The two largest clusters (99 and 83 equations) centered on Maxwell’s equations, confirming the internal consistency of electromagnetic theory.
+The system identified complex relationships between the wave equation for electromagnetic fields and Maxwell’s equations through multiple substitutions, though the resulting expressions contain parsing artifacts that require further investigation.
+These clusters validate the framework’s ability to recognize the mathematical coherence of established physical theories.
+4.2.3 Electromagnetic-Fluid Coupling (Cluster #8)
+Cluster #8 revealed an unexpected synthesis between fluid dynamics and electromagnetism, identified by an earlier version of the code that still had parsing issues but nonetheless uncovered this intriguing connection (full report in Supplementary Materials).
+Starting with Reynolds number as backbone:
+$Re=\frac{\rho vL}{\eta}$ (5)
+The system applied the following substitutions from electromagnetic equations in the cluster:
+  • E=Bv+F/qE=-Bv+F/q (from Lorentz Force (Complete))
+  • $t=I\epsilon/L$ (from Inductance EMF)
+  • $v=\epsilon/(BL)$ (from Motional EMF)
+  • $t=-d\epsilon/d\Phi_{B}$ (from Faraday’s Law)
+  • $t=-\epsilon/\Phi$ (from Lenz’s Law Direction)
+This produced the simplified expression:
+$\frac{\epsilon\rho}{B\eta}$ (6)
+This result represents a dimensionless parameter coupling electromagnetic and fluid properties, analogous to the Magnetic Reynolds Number in magnetohydrodynamics, demonstrating how the framework can identify cross-domain mathematical structures even without understanding the underlying physics.
+4.3 Error Detection as Knowledge Auditing
+4.3.1 Dimensional Catastrophe (Cluster #4)
+The four-momentum cluster exposed an error in the knowledge base. Starting with the relativistic invariant as backbone:
+$\mu pp_{\mu}=c^{2}mr^{2}$ (7)
+The system applied the following substitutions from equations in the cluster:
+  • $p=E/c-cm^{2}$ (from Four-Momentum - Relativistic Invariant)
+  • $m=\sqrt{E-cp}/c$ (from Four-Momentum - Relativistic Invariant)
+  • $c=m_{e}$ (from Compton Scattering - Wavelength Shift)
+This produced:
+$\mu pp_{\mu}=m_{e}r^{2}\sqrt{E-m_{e}p}$ (8)
+With residual: $m_{e}r^{2}\sqrt{E-m_{e}p}+\mu pp_{\mu}$
+The critical error is the substitution $c=m_{e}$, which is dimensionally incorrect—equating velocity [L/T] with mass [M].
+This error originates from a mis-parsed Compton scattering equation where the system incorrectly extracted a relationship between the speed of light and electron mass, likely from the Compton wavelength formula $\lambda_{C}=h/(m_{e}c)$ where $m_{e}$ and $c$ appear together but represent fundamentally different physical quantities.
+4.3.2 Category Confusion (Cluster #16)
+The framework combined Bragg’s Law for crystal diffraction:
+$2d\sin\theta=n\lambda$ (9)
+with substitutions from interference equations:
+  • $m=d\sin\theta/\lambda$ (from Interference - Young’s Double Slit)
+  • $a=m\lambda/\sin\theta$ (from Single Slit Diffraction (Minima))
+  • $m=(d\sin\theta-\lambda/2)/\lambda$ (from Interference - Young’s Double Slit (dsi))
+This produced the tautology True, but mixed different physical contexts—crystalline solids versus free space.
+4.4 The Unexpected Value of Errors: Analog Gravity (Cluster #15)
+Cluster #15 combined 11 equations from fluid dynamics and electromagnetism. Starting with Bernoulli’s equation:
+$P+\frac{1}{2}\rho v^{2}+\rho gh=\text{constant}$ (10)
+The system applied substitutions:
+  • $h=v/(gr^{2})$ (from Torricelli’s Law)
+  • $g=v/(hr^{2})$ (from Torricelli’s Law)
+  • $c=(\mu_{r}-1)/\chi_{m}$ (from Magnetic Susceptibility)
+  • $P=P_{0}+\rho gh$ (from Hydrostatic Pressure)
+This produced:
+$\frac{(\mu_{r}-1)^{2}}{h^{2}m^{2}}$ (11)
+This mixes magnetic permeability with fluid dynamical variables—dimensionally inconsistent and physically incorrect.
+However, this error points to analog gravity research, where:
+  • Sound waves in fluids obey equations mathematically identical to quantum fields in curved spacetime
+  • Fluid velocity fields create effective metrics analogous to gravitational geometry
+  • Sonic horizons mirror black hole event horizons
+  • Navier-Stokes equations in $(p+1)$ dimensions correspond to Einstein equations in $(p+2)$ dimensions
+Near the horizon limit, Einstein’s equations reduce to Navier-Stokes, suggesting gravity might emerge from microscopic degrees of freedom.
+4.5 Statistical Summary
+Table 3: Cluster Analysis Results
+Category Count Examples
+Theory Validation 2 Klein-Gordon/Dirac, Maxwell consistency
+Novel Synthesis 8 Transport phenomena, Reynolds-EMF
+Tautologies 5 Planck’s Law circular substitutions
+Dimensional Errors 2 $c=m_{e}$, knowledge base errors
+Category Errors 3 Bragg/Young confusion
+Provocative Failures 4 Analog gravity connection
+Insufficient Data 6 No valid substitutions
+Total 30 80% produced interpretable results
+The framework operates at the syntactic level, recognizing mathematical patterns without understanding physical causality.
+This limitation, combined with parsing issues in early iterations, requires human interpretation to distinguish profound connections from tautologies or errors.
+However, even computational failures can suggest legitimate research directions, as demonstrated by the analog gravity connection.
 5 Discussion
-The results may support my hypothesis that a graph neural network, when combined with a methodologically sound graph representation, can identify statistically significant mathematical patterns in a corpus of physics equations.
-The model’s ability to achieve a robust and high AUC score ($0.9742\pm 0.0030$) after a semantic disambiguation and validation process indicates the discovered relationships may reflect true structural similarities in the mathematical formalism, not notational artifacts or model overfitting.
-The discovered connections must be interpreted as syntactic analogies, not as direct physical or causal relationships.
-For instance, the link between Photon Energy (Quantum Mechanics) and the relativistic Four-Momentum invariant (Relativity) does not imply a new physical theory connecting them.
-Rather, it suggests a “meta-law”: a shared mathematical grammar used to describe critical phenomena, even when they occur in vastly different energy regimes.
-The model is “semantically blind”— it recognizes syntax, not physical causality. Therefore, its predictions must be interpreted as hypotheses that require further theoretical investigation.
+This proof of concept demonstrates a framework capable of systematic pattern discovery in physics equations.
+The framework functions as a computational lens revealing hidden structures and inconsistencies difficult to discern at human scale.
+It does not replace physicists, of course, but fully developed could augment their capabilities through systematic exploration of mathematical possibility space.
+The framework’s value lies not in autonomous discovery but in its role as a computational companion—a tireless explorer of mathematical possibility space that surfaces patterns, errors, and unexpected connections for human interpretation.
+Even its failures are productive, transforming the vast combinatorial space of physics equations into a curated set of computational hypotheses worthy of expert attention.
+The transformation of mathematical pattern into physical understanding remains fundamentally human.
+But by automating the pattern-finding, the framework frees physicists to focus on what they do best: interpreting meaning, recognizing significance, and building the conceptual bridges that transform equations into understanding of nature.
 6 Conclusion and Future Work
-I have developed and preliminarily validated a GNN-based framework capable of mapping the mathematical structure of physical law.
-Our proposed GAT model achieves state-of-the-art performance (AUC = 0.9742) on this novel task, demonstrably outperforming classical and established GNN baselines in a statistically-validated comparison.
-By addressing key methodological challenges—from semantic disambiguation to objective edge weight definition—we have shown that this approach can identify non-trivial, physically interesting, and statistically significant patterns.
+In this preliminary work, I have developed and tested a GNN-based framework capable of mapping the mathematical structure of physical law, and acting as a computational auditor.
+My GAT model achieves high performance on this novel task, and the primary contribution of this work lies in the subsequent symbolic analysis.
+This analysis suggests the framework has a potential multi-layered capacity to: (i) verify the internal consistency of foundational theories, (ii) help debug knowledge bases by identifying errors and tautologies, (iii) synthesize mathematical structures analogous to complex physical principles, and (iv) provide creative provocations from its own systemic failures.
 6.1 Current Limitations and Future Directions
-  • Database: The immediate next step is to expand the equation database to over 1,000 equations.
-This richer corpus will not only broaden the map but also enable the encoding of deeper mathematical structures—such as the presence of specific operators, derivatives, or symmetries—as features, moving the analysis from a syntactic to a more profound structural level.
-  • Dynamic Methodological Refinements: Future work will explore more adaptive methodologies.
-This includes developing learnable edge-weighting schemes that go beyond a fixed linear combination and extending our disambiguation protocol to act as a general-purpose “auditor” for standardizing notational conventions across different scientific knowledge bases.
-  • Advanced Statistical Analysis: Building on the rigorous FDR validation presented, future research can focus on more nuanced metrics that assess the novelty and non-triviality of discoveries.
-We also plan to explore methods from causal inference to investigate the potential directionality of the identified mathematical relationships.
-  • Interactive Tools for Expert Validation: To bridge the gap between computational patterns and physical insight, we plan to develop interactive visualization tools that facilitate an expert-in-the-loop process to validate and interpret the most promising machine-generated analogies.
+This work, while promising, represents an initial step. The path forward is clear and focuses on several key areas:
+  • Systematic Parameter Exploration: The framework requires systematic testing with different hyperparameter configurations to identify optimal settings for various physics domains.
+Different weight combinations in the importance scoring and clustering algorithms may reveal distinct classes of mathematical relationships, generating a vast number of hypotheses that require careful evaluation.
+  • AI-Assisted Hypothesis Screening: The system currently generates hundreds of potential cross-domain connections, many of which are spurious or trivial.
+Future work should integrate large language models or other AI systems to perform initial screening of these computational hypotheses, filtering out obvious errors, tautologies, and dimensionally inconsistent results before human expert review.
+This would create a multi-stage pipeline: graph-based generation, AI screening, and expert validation.
+  • Database Expansion: The immediate next step is to expand the equation database.
+A richer and broader corpus would enable the encoding of deeper mathematical structures, moving the analysis from a syntactic to a more profound structural level.
+  • Generalization as a Scientific Auditor: Future work will focus on generalizing the framework beyond physics to other formal sciences.
+This includes refining the disambiguation protocol to act as a general-purpose “auditor” for standardizing notational conventions across different scientific knowledge bases.
+  • Collaboration with Domain Experts: To bridge the gap between computational patterns and physical insight, future work must involve an expert-in-the-loop process.
+Collaboration with theoretical physicists is essential to validate, interpret, and build upon the most promising machine-generated analogies and audit reports.
 6.2 Broader Implications
-This work opens several avenues for the broader scientific community, including applications in education, philosophy of science, and other scientific domains like chemistry and biology.
+This work may open several avenues for the broader scientific community.
+As an auditing tool, it could potentially be used to systematically check the consistency of large-scale theories.
+As an educational tool, it might help students visualize the deep structural connections that unify different areas of science.
+More broadly, this research contributes to the emerging field of computational epistemology, developing methods to study the structure and coherence of scientific knowledge.
+Ultimately, this framework is presented as a tangible step toward a new synergy between human intuition and machine computation, where AI may serve as a tool to augment and stimulate the quest for scientific understanding.
 7 Data and Code Availability
-All code, the cleaned dataset, and model weights will be available at: https://github.com/kingelanci/graphysics.
+All code, cleaned dataset, and model weights available at: https://github.com/kingelanci/graphysics.
+8 Supplementary Materials
+Complete prediction distributions and analysis results available at the GitHub repository:
+  • Full distribution of AUC ¿ 0.92 prediction scores
+  • Complete symbolic analysis for all 30 clusters
+  • Bootstrap validation logs
+  • All generated cross-domain hypotheses (not just selected examples)
 Correspondence
-Correspondence should be addressed to Massimiliano Romiti (massimiliano.romiti@acm.org).
+Correspondence: Massimiliano Romiti (massimiliano.romiti@acm.org).
 Acknowledgments
-The author thanks the arXiv endorsers and the anonymous reviewers for their valuable feedback. Special gratitude to the ACM community for professional development resources.
+The author thanks the arXiv endorsers for their support.
+Special gratitude to the ACM community for professional development resources and to online physics communities for discussions that helped clarify physical interpretations.
+While AI tools were employed for auxiliary tasks such as code debugging, literature search, formatting assistance, and initial draft organization, all scientific content, analysis, interpretations, and conclusions are solely the work of the author.
+Any errors or misinterpretations remain the author’s responsibility.